OpenStudy (anonymous):

alright i dont know how to do this. x=16-4x 3x+4y=8

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

first make x at left had side x+4x= 16 5x = 16 x = 16/5 then 3x+4y =8 4y =8 -3x y = 1/4(8 - 3x) insert x y = 1/4(8 - 3(16/5)) y =-2/5

OpenStudy (anonymous):

but my teacher says it shouldnt be a fraction

OpenStudy (anonymous):

you type the question wrong

OpenStudy (anonymous):

well thats how it is on my paper

OpenStudy (anonymous):

for : x=16-4y; 3x+4y=8 x =5 and y = 11/4

OpenStudy (anonymous):

okay thanks:)

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

like wrok it out please?

OpenStudy (anonymous):

efgvbknlmkjhgkmdlsavcgzmsbtrgftiykhejbl45tiuygftidjbwqlhe;trdsakewjrtyuiutr

OpenStudy (anonymous):

x=16-4y as first equation, 3x+4y=8 change to 3x=8-4y then to x=1/3(8-4y) and take as second equation first equation = second equation as x = x 16-4y =1/3(8-4y) 3(16-4y)= 8-4y 48-12y = 8 -4y 48-8 = 12y -4y 40 = 8y 5 = y then insert y=5 into x=16-4y x =16-4(5) x= -4 then x=-4 and y=5

OpenStudy (anonymous):

sorry for miscalculating b4

OpenStudy (anonymous):

its okay i was just confused